OpenStudy (gabylovesyou):

Lance can paint the room 5 hours, while Kim can paint the same room in 4 hours. How many hours would it take for them to paint the room together? 9 20 twenty ninths nine twentieths

OpenStudy (ryuga):

twenty ninths

OpenStudy (welshfella):

find out what farction of the room they can paint in 1 hour in terms of time

OpenStudy (crystal_bliss):

nine twentieths. you just add them and then turn it into a number

OpenStudy (welshfella):

lance paints 1/5 of room in 1 hour, Kim paints 1/4 Working together they paint 1 / t so 1/5 + 1/4 = 1/t solve for t

OpenStudy (welshfella):

I wish people wouldn't just give the answer

OpenStudy (gabylovesyou):

yea i get 9/20.. thank you.

OpenStudy (welshfella):

its not 9/20 its 20 / 9 multiply each fraction by the LCM 20t, we get 4t + 5t = 20 t = 20/9

OpenStudy (gabylovesyou):

? o-o

OpenStudy (gabylovesyou):

OHHHHHh i seeee ok... nvm.

OpenStudy (welshfella):

lol

OpenStudy (gabylovesyou):

lol cx thanks.

OpenStudy (welshfella):

yw